OpenStudy (anonymous):

Solve 3x - 4 ≤ 2 or 2x + 11 ≥ -1. all real?

HanAkoSolo (jamierox4ev3r):

3x-4\(\le\)2 or 2x+11\(\ge\)-1 +4+4 -11 -11 ------------------------------ 3x\(\le\)6 or 2x\(\ge\)-12 /3 /3 /2 /2 -------------- x\(\le\)2 or x\(\ge\)-6

HanAkoSolo (jamierox4ev3r):

lol that took me a while :P

HanAkoSolo (jamierox4ev3r):

do you understand @TheGstudent23 ?

OpenStudy (anonymous):

oh ok so its {x | -6 ≤ x ≤ 2}?

HanAkoSolo (jamierox4ev3r):

yep thats right :)

OpenStudy (anonymous):

ok thank you